OpenStudy (anonymous):

nput the equation of the given line in standard form. The line with m = 2/3 and passing through (1, 1)

OpenStudy (anonymous):

(x1,y1)=(1,1) y-y1=m*(x-x1) y-1=2(x-1)/3 3y-3=2x-2 2x-3y+1=0